OpenStudy (anonymous):

what is the definite integral of (1+x)^n when n is not equal to -1

OpenStudy (raden):

u need integrate by u-sub, first

OpenStudy (raden):

let u = 1+x du = dx

OpenStudy (raden):

now, that integral can be int u^n du = 1/(n+1) u^(n+1) + c

OpenStudy (raden):

subtitute back that u = 1+x

OpenStudy (raden):

so, it can be